OpenStudy (anonymous):

Find the surface area of a sphere that has a volume of 288π cu in.

OpenStudy (anonymous):

answer choices are 48 ,144 and 20,736

OpenStudy (anonymous):

A=4*pi*r^2 A=4*36*pi =144*pi

OpenStudy (anonymous):

you were right zed.

OpenStudy (anonymous):

omar was right, just a little calculator error. He did all the hard work :)

OpenStudy (anonymous):

ok thank you both

OpenStudy (anonymous):

volume V=(4/3)*pi*r^3 r: radius.....r=6 in surface area A=4*pi*r^2 A=144pi sq in Thanx guys

OpenStudy (anonymous):

Volume of a sphere is given by the formula (4/3)*pi*r^3 Surface area of a sphere is given by the formula 4*pi*r^2 We need to find r 4/3 * pi * r^3 = 288 * pi So 4/3 * r^3 = 288 and r = cuberoot (3/4 * 288) = cuberoot (216) = 6 So the surface area of the sphere would be 4 * pi * 6^2 = 144 * pi = 452.389 cu in